# Data Augmentation Through Random Style Replacement

### v1

@@ I Introduction
Recent advancements in deep learning have driven significant progress in a wide range of computer vision tasks, including image classification, object detection, and semantic segmentation [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. Despite these advances, many of these tasks continue to face a fundamental bottleneck: a lack of sufficient labeled data [12, 13, 14, 15, 16]. Annotating large-scale datasets is both time-consuming and costly, which can limit the applicability of deep neural networks in specialized or rapidly evolving domains. To mitigate this issue, data augmentation techniques are heavily utilized, artificially expanding and diversifying the training set so that models generalize more effectively.
@@
Data augmentation has garnered significant attention in supervised learning research across a wide range of domains—including computer vision [17, 18, 19, 20], natural language processing [21, 22, 23, 24], graph learning [25, 26, 27, 28, 29, 30], time series prediction [31, 32], large language model [33], and the internet of things [34, 35] — due to its ability to increase both the volume and diversity of training data, thereby enhancing model generalization and mitigating overfitting. Broadly, data augmentation strategies can be grouped into two categories: generative methods, which utilize models like Variational Autoencoders (VAEs), Generative Adversarial Networks (GANs), Large Language Models (LLMs), or diffusion-based frameworks to synthesize new data [36, 37, 38, 39, 40, 41, 42, 43]; and traditional methods, which rely on transformations such as random cropping, flipping, rotations, color jittering, and histogram equalization to modify existing samples. While both approaches aim to expose the model to a wider variety of conditions, thus reducing overfitting, traditional augmentation strategies may not fully capture the complexity of real-world variability. Consequently, several studies have explored more refined methods, including style augmentation and random erasing [44, 45, 46, 47, 48]. Style augmentation employs style transfer to alter the visual attributes of training images while preserving their semantic content, thereby increasing robustness to differences in texture, color, and contrast. Random erasing, on the other hand, randomly occludes or replaces subregions of an image, making models more resilient to missing or corrupted information. In this paper, we revisit these traditional approaches—particularly focusing on their potential to advance the efficacy of data augmentation in supervised learning.
@@ VI Future Work
Our random style replacement method has shown promising results, yet further validation is needed to confirm its wider applicability. It is crucial to test this technique across various datasets and tasks to establish its generalizability and identify any limitations. Additionally, the convergence speed observations require confirmation through further experiments involving diverse datasets and network architectures. Moreover, integrating Large Language Models (LLMs) guided approaches [52, 53, 54, 55, 56, 57, 58, 59] could enhance the method. These approaches would use LLMs to guide style replacement, potentially selecting optimal subregions for style transfer based on foreground and background information, thus enabling more meaningful and effective transformations.

### v2

@@ I Introduction
Recent advancements in deep learning have driven significant progress in a wide range of computer vision tasks, including image classification, object detection, and semantic segmentation [1, 2]. Despite these advances, many of these tasks continue to face a fundamental bottleneck: a lack of sufficient labeled data [3, 4]. Annotating large-scale datasets is both time-consuming and costly, which can limit the applicability of deep neural networks in specialized or rapidly evolving domains. To mitigate this issue, data augmentation techniques are heavily utilized, artificially expanding and diversifying the training set so that models generalize more effectively.
@@
Data augmentation [5] has garnered significant attention in supervised learning research across a wide range of domains—including computer vision [6, 7], natural language processing [8], graph learning [9, 10], and large language model [11, 12] — due to its ability to increase both the volume and diversity of training data, thereby enhancing model generalization and mitigating overfitting. Broadly, data augmentation strategies can be grouped into two categories: generative methods, which utilize models like Variational Autoencoders (VAEs) [13], Generative Adversarial Networks (GANs) [14], Large Language Models (LLMs), or diffusion-based frameworks [15] to synthesize new data; and traditional methods, which rely on transformations such as random cropping, flipping, rotations, color jittering, and histogram equalization to modify existing samples. While both approaches aim to expose the model to a wider variety of conditions, thus reducing overfitting, traditional augmentation strategies may not fully capture the complexity of real-world variability. Consequently, several studies have explored more refined methods, including style augmentation and random erasing [16, 17, 18]. Style augmentation employs style transfer [19] to alter the visual attributes of training images while preserving their semantic content, thereby increasing robustness to differences in texture, color, and contrast. Random erasing, on the other hand, randomly occludes or replaces subregions of an image, making models more resilient to missing or corrupted information. In this paper, we revisit these traditional approaches—particularly focusing on their potential to advance the efficacy of data augmentation in supervised learning.
@@ VI Future Work
Our random style replacement method has shown promising results, yet further validation is needed to confirm its wider applicability. It is crucial to test this technique across various datasets and tasks to establish its generalizability and identify any limitations. Additionally, the convergence speed observations require confirmation through further experiments involving diverse datasets and network architectures. Moreover, integrating Large Language Models (LLMs) guided approaches [23] could enhance the method. These approaches would use LLMs to guide style replacement, potentially selecting optimal subregions for style transfer based on foreground and background information, thus enabling more meaningful and effective transformations.